The product of two consecutive integers is 4 less than 4 times their sum. Find them
Answer by stanbon(75887)

The product of two consecutive integers is 4 less than 4 times their sum.
1st: x
2nd: x+1
-----
Equation:
x(x+1) = = 4(2x+1)+4
x^2+x = 8x+8
x(x+1) = 8(x+1)
x = 8
x+1 = 9
===============
Cheers,
Stan H.